OpenStudy (gabbyalicorn):

What is the approximate circumference of a circle with a radius of 30 inches? Use π ≈ 3.14. A. 94.2 inches B. 62.8 inches C. 188.4 inches D. 47.1 inches

OpenStudy (gabbyalicorn):

so what im doing is 3.14 x 30 = 94.2 / 2 47.1 can someone double check if this is correct please.

OpenStudy (gabbyalicorn):

@pinklion23 @iGreen

OpenStudy (spring98):

it is correct i got the same answer: )

OpenStudy (gabbyalicorn):

Oh, hazaa!

OpenStudy (spring98):

lol

OpenStudy (igreen):

No, that's incorrect

OpenStudy (igreen):

The formula is \(\sf 2 \pi r\), not \(\sf \dfrac{\pi r}{2}\).

OpenStudy (gabbyalicorn):

so 2 x 3.14 x 30 ?

OpenStudy (igreen):

Yep

OpenStudy (gabbyalicorn):

which would be... 188.4

OpenStudy (spring98):

opps thats what i ment!

OpenStudy (deana99):

So it will be C.

OpenStudy (igreen):

Yep, that's our answer.

OpenStudy (gabbyalicorn):

Thanks, Igreen!

OpenStudy (igreen):

Np